# How to Collaborate: Towards Maximizing the Generalization Performance in Cross-Silo Federated Learning

### v1

@@ Abstract
Federated learning (FL) has attracted vivid attention as a privacy-preserving distributed learning framework. In this work, we focus on cross-silo FL, where clients become the model owners after training and are only concerned about the model’s generalization performance on their local data. Due to the data heterogeneity issue, asking all the clients to join a single FL training process may result in model performance degradation. To investigate the effectiveness of collaboration, we first derive a generalization bound for each client when collaborating with others or when training independently. We show that the generalization performance of a client can be improved only by collaborating with other clients that have more training data and similar data distribution. Our analysis allows us to formulate a client utility maximization problem by partitioning clients into multiple collaborating groups. A hierarchical clustering-based collaborative training (HCCT) scheme is then proposed, which does not need to fix in advance the number of groups. We further analyze the convergence of HCCT for general non-convex loss functions which unveils the effect of data similarity among clients. Extensive simulations show that HCCT achieves better generalization performance than baseline schemes, whereas it degenerates to independent training and conventional FL in specific scenarios.
@@ 1 Introduction
There have been a few attempts to improve the utility of collaboration in FL. A recent work [12] proposed to only incorporate the clients that find the global model appealing in FL while allowing others to train independently. Nevertheless, those independent clients, which have data distributions different from most clients, may not enjoy satisfactory performance by making a binary decision of whether to participate in FL or not. To remedy the deficiency of a single global model, some works [13, 14, 15] introduce multiple global models to FL system to the FL system. Specifically, they cluster all clients into several groups, and the clients in the same group cooperatively train a global model. The clustering criterion is tailored to minimize the training error of clients, which, however, is not sufficient to guarantee good performance on unseen or new data [16]. To enhance clients’ generalization performance, the design of an appropriate metric for partitioning clients remains an unresolved challenge. Besides, determining the number of groups without additional knowledge of local data distribution is a notoriously challenging problem. As a more general consideration, some studies resort to game theory to formulate the client partition problem as a hedonic game without fixing the number of groups [17, 18, 19, 20, 21]. However, these works focus on the theoretical aspects of this game, e.g., stability and equilibrium, and they do not provide any practical or efficient algorithm to solve the problem.
@@
  • By collaborating, clients can benefit from more training data, but may also suffer from data heterogeneity. To study in detail this phenomenon, we analyze the local test error of a client when it joins any group or trains a model independently. Our analytical results unveil that the generalization performance of a client can be improved by involving more training data samples yet excluding the collaborators with diverged data distribution.
  • Motivated by this analysis, we define the utility of each client as an upper bound of the test error, where the divergence of data distribution is approximated by the gradient distance. Then, we formulate a client utility maximization problem by designing the collaboration among clients. To solve the problem efficiently, we propose a hierarchical clustering-based collaborative training scheme, coined HCCT, which identifies the proper collaboration patterns for clients during training. It is worth noting that the number of groups in HCCT is automatically determined without requiring additional tuning. Considering the computation complexity of evaluating client similarity, we further propose an efficient implementation of HCCT.
@@ 2 Related Works
Fair collaboration in single-model FL. Considering the client requirements, Cho et al. [12] proposed a training scheme named MAXFL, where each client participates in FL only if it finds the global model appealing, namely, if the global model yields a smaller training loss than a self-defined threshold. This threshold is defined as the difference between the current global model and a local model optimized in an independent training process, which, however, introduces additional computation cost. More importantly, the binary decision used in this work excludes the possibility that those independent clients may formulate a group to further improve their generalization performance. Another stream of work [25, 26] aims to achieve fair model performance among clients by modifying the global training objective. For example, [25] proposed to give a higher weight to the model updates from clients with lower training accuracy. Nevertheless, these attempts still focus on the setup with a single global model, which limits the performance improvement for most clients.
Clustered federated learning. To satisfy different local objectives, clustered FL generates multiple global models by clustering clients into several groups [13, 14, 15]. For example, an iterative clustering algorithm for FL named IFCA was proposed in [13], where client group identity estimation and federated model training are carried out alternatively. To estimate the group identity, clients are required to evaluate each global model on their local data to find the one achieving the minimal training loss. However, this process incurs additional training cost that increases linearly with the number of groups. More importantly, the clustering criterion in these works emphasizes the training loss but fails to reflect the generalization performance on the unseen data. Furthermore, selecting the number of groups is known to be challenging as the server does not know local data distribution.
@@ 3.1 Cross-Silo FL
Here $l(h(\mathbf{x};\mathbf{w}),y)$ is the loss function (e.g., the categorical cross-entropy) computed between the predicted label $h(\mathbf{x};\mathbf{w})$ and the ground-truth label $y$, and $\mathcal{P}_{i}^{te}$ denotes the distribution of test data. Although the test data is typically unknown during training, it is commonly assumed that the training data and test data are generated from the same underlying distribution [31], i.e.,
@@
In real-world scenarios, clients have various data distributions and different training dataset sizes [11, 32]. To be specific, some clients may own a limited number of data samples and thus seek assistance from others for collaboration to improve the model performance. To describe this scenario, without loss of generality, we assume that clients are partitioned into $K$ groups with indices of $k\in[K]$. Note that the number of groups $K$ is a priori unknown. Each group consists of a set of clients $\mathcal{C}_{k}\subset\mathcal{N}$, and a client belongs to only one group. With the available training data $\hat{\mathcal{D}}_{k}^{tr}\triangleq\cup_{i\in\mathcal{C}_{k}}\mathcal{D}_{i}^{tr}$, clients in group $k$ train a model by minimizing the following objective:
@@
In the training epoch $t$, the training process of each group is detailed as follows. Each client $i$ receives the global model $\hat{\mathbf{w}}_{G_{i}^{t}}^{t}$ from its corresponding group $G_{i}^{t}\in[K]$ and optimizes this model by using mini-batch stochastic gradient descent (SGD) for $Q$ steps. Specifically, at step $q\in[Q]$ of epoch $t$, client $i$ updates the local model according to
@@
$\mathbf{g}_{i}^{t}\triangleq\sum_{q=1}^{Q}\sum_{(\mathbf{x},y)\in\mathcal{B}_{i}^{t,q}}\nabla_{\mathbf{w}}l(h((\mathbf{x};\mathbf{w}_{i}^{t,q}),y),$ (5)
@@
It is worth noting that clients have various data distribution [1, 4], and thus some clients’ data may differ significantly from others. However, under strict privacy protection, the local data distribution is only visible to the client itself [33]. If the process of joining a group is performed arbitrarily, clients may suffer from performance degradation due to the data heterogeneity issue [11]. To clearly demonstrate this phenomenon, we present an example in the next subsection.
@@ 3.2 Motivating Example
We consider a cross-silo FL system with $N=3$ clients. They collaborate to classify data samples extracted from the CIFAR-10 [34] dataset by training a convolutional neural network (CNN) model with four convolutional layers. We assume that clients $1$ and $2$ have the same data distribution and client $3$ has different data distribution deviating from the two others. To be concrete, clients $1$ and $2$ have respectively $20\%$ and $80\%$ of five classes of training data, while client $3$ has all the training data samples of the other five classes. We simulate different collaboration patterns among three clients and evaluate their test errors, as shown in Table II, where the best generalization performances are highlighted in bold while the worst ones are underlined.
@@
$*{}^{*}$ $(1,2)$ $\mathbf{23.54\pm 1.40}$ $\mathbf{25.28\pm 0.64}$ $\mathbf{15.90\pm 1.19}$ $\mathbf{21.58}$ $\mathbf{4.10}$
@@
$*{}^{*}$wo clients collaborate on the global model, while the other client trains independently.
@@ 4.1 Theoretical Analysis of Generalization Performance
Consider the loss function with $\mu$-strong convexity and $L$-Lipschitz continuity. For a prediction function $h$ and any $\delta>0$, the following holds:
@@
with probability $1-\delta$, where $d_{1}\left(\hat{\mathcal{P}}_{G_{i}},\mathcal{P}_{i}\right)$ is the distribution divergence (e.g., $L^{1}$-divergence or $\mathcal{H}$-divergence) between $\hat{\mathcal{P}}_{G_{i}}$ and $\mathcal{P}_{i}$, and $\lambda=\min\left\{\mathbb{E}_{\mathcal{P}_{i}}\left[f_{G_{i}}(\mathbf{x})-f_{i}(\mathbf{x})\right],\mathbb{E}_{\hat{\mathcal{P}}_{G_{i}}}\left[f_{G_{i}}(\mathbf{x})-f_{i}(\mathbf{x})\right]\right\}$ with true labeling functions $f_{i}(\cdot)$ and $f_{G_{i}}(\cdot)$.
@@
Please refer to Appendix A.
@@
The above results show that any client should seek collaborators with similar training data distribution and more data samples. However, the data distribution of clients cannot be disclosed to others due to privacy concerns. To solve this problem, in the following subsection, we propose a method to estimate the data divergence and formulate the client utility maximization problem.
4.2 Problem Formulation
@@
$s\left(i,G_{i}\right)=\cos\langle\mathbf{g}_{i}^{t},\mathbf{g}_{G_{i}}^{t}\rangle=\frac{\left\langle\mathbf{g}_{i}^{t},\mathbf{g}_{G_{i}}^{t}\right\rangle}{\left\|\mathbf{g}_{i}^{t}\right\|_{2}\left\|\mathbf{g}_{G_{i}}^{t}\right\|_{2}}.$ (8)
@@
$U_{i}(G_{i})\triangleq-\frac{\alpha}{\hat{D}_{G_{i}}}+s\left(i,G_{i}\right)+\beta,$ (9)
where $\alpha>0$ is a trade-off constant, and $\beta>0$ is a large constant ensuring non-negative utility without impacting system design. We note that $\alpha$ balances between training data volume and gradient similarity. If clients have sufficient data, they prefer collaborating with similar clients by adopting a small $\alpha$; otherwise, they are hungry for more data and choose a large $\alpha$. The impact of $\alpha$ is discussed further in Section 7.4 via simulations. According to the definition in (9), a client can improve its utility by involving more training data and increasing the similarity between gradients.
@@
Finding an optimal solution for Problem (P1) is $\mathcal{NP}$-hard in general [20]. One straightforward approach is to search over all $B_{N}$ possible partitions, where $B_{N}$ denotes the Bell number of a set with $N$ elements. This approach, however, introduces extremely high computation complexity. We notice that if $\alpha=0$, Problem (P1) reduces to a $k$-means problem which partitions $N$ clients into $K$ groups to minimize the within-cluster variance, i.e., the distance between each point $\mathbf{g}_{i}^{t}$ and its corresponding center $\mathbf{g}_{G_{i}}^{t}$. There are many classic methods for solving this problem, e.g., $k$-means clustering [37], $k$-means++ [38] and hierarchical clustering [39, 40] algorithms. However, the presence of the constraint $\alpha>0$ makes it unfeasible to define a suitable distance metric to solve Problem (P1). Consequently, these methods cannot be directly employed. Drawing inspiration from the aforementioned clustering algorithms, we introduce an algorithm to efficiently address the client partition in Problem (P1) in the following section.
@@ 5 Hierarchical Clustering-based Collaborative Training Scheme
[…]
@@
where $\hat{D}_{k^{\prime}}=\sum_{j\in\mathcal{C}_{k^{\prime}}}D_{j}$ is the number of training data samples of clients in $\mathcal{C}_{k^{\prime}}$. For these clients in group $k^{\prime}$, we recompute their utilities and sum them up as the utility of this group, i.e,
@@
It is worth noting that there is no need to select beforehand the number of groups $K$. In every partitioning step, the value of $K$ will be deducted by one automatically. Thus, the number of groups is implicitly determined by the stopping criterion. Specifically, the process of client partition is terminated if there is only one group left, i.e., $K=1$, or there is no benefit in forming collaborations between any two groups, i.e.,
@@
The process of client partition is summarized in Algorithm 2, and we show an example with six clients in Fig. 2.

### v2

@@ Abstract
Federated learning (FL) has attracted vivid attention as a privacy-preserving distributed learning framework. In this work, we focus on cross-silo FL, where clients become the model owners after training and are only concerned about the model’s generalization performance on their local data. Due to the data heterogeneity issue, asking all the clients to join a single FL training process may result in model performance degradation. To investigate the effectiveness of collaboration, we first derive a generalization bound for each client when collaborating with others or when training independently. We show that the generalization performance of a client can be improved by collaborating with other clients that have more training data and similar data distributions. Our analysis allows us to formulate a client utility maximization problem by partitioning clients into multiple collaborating groups. A hierarchical clustering-based collaborative training (HCCT) scheme is then proposed, which does not need to fix in advance the number of groups. We further analyze the convergence of HCCT for general non-convex loss functions which unveils the effect of data similarity among clients. Extensive simulations show that HCCT achieves better generalization performance than baseline schemes, whereas it degenerates to independent training and conventional FL in specific scenarios.
@@ 1 Introduction
There have been a few attempts to improve the utility of collaboration in FL. A recent work [12] proposed to only incorporate the clients that find the global model beneficial in collaborative training while allowing others to train independently. Nevertheless, those independent clients, which have different data distributions from most clients, may not enjoy satisfactory performance by making a binary decision of whether to participate in FL or not. To remedy the deficiency of a single global model, some works [13, 14, 15] introduce multiple global models to FL system. Specifically, they cluster all clients into several groups, and the clients in the same group cooperatively train a global model. The clustering criterion is tailored to minimize the training error of clients, which, however, is not sufficient to guarantee good performance on unseen or new data [16]. To enhance clients’ generalization performance, the design of an appropriate metric for partitioning clients remains an unresolved challenge. Besides, determining the number of groups without additional knowledge of local data distribution is a notoriously challenging problem. As a more general consideration, some studies resort to game theory to formulate the client partition problem as a hedonic game without fixing the number of groups [17, 18, 19, 20, 21]. However, these works focus on the theoretical aspects of this game, e.g., stability and equilibrium, and they do not provide any practical and efficient algorithm to solve the problem.
@@
  • By collaborating, clients can benefit from more training data, but may also suffer from data heterogeneity. To study in detail this phenomenon, we analyze the local test error of a client when it joins any group or trains a model independently. Our analytical results unveil that the generalization performance of a client can be improved by involving more training data samples yet excluding the collaborators with diverged data distributions.
  • Motivated by this analysis, we define the utility of each client as an upper bound of the test error, where the divergence of data distribution is approximated by the gradient distance. Then, we formulate a client utility maximization problem by designing the collaboration among clients. To solve the problem efficiently, we propose a hierarchical clustering-based collaborative training scheme, coined HCCT, which identifies the proper collaboration patterns for clients during training. It is worth noting that the number of groups in HCCT is automatically determined without requiring additional tuning. Considering the computational complexity of evaluating client similarity, we further propose an efficient implementation of HCCT.
@@ 2 Related Works
Fair collaboration in single-model FL. Considering the client requirements, Cho et al. [12] proposed a training scheme named MAXFL, where each client participates in FL only if it finds the global model appealing, namely, if the global model yields a smaller training loss than a self-defined threshold. This threshold is defined as the difference between the current global model and a local model optimized in an independent training process, which, however, introduces additional computation costs. More importantly, the binary decision used in this work excludes the possibility that those independent clients may formulate a group to further improve their generalization performance. Another stream of work [25, 26] aims to achieve fair model performance among clients by modifying the global training objective. For example, [25] proposed to give a higher weight to the model updates from clients with lower training accuracy. Nevertheless, these attempts still focus on the setup with a single global model, which limits the performance improvement for most clients.
Clustered federated learning. To satisfy different local objectives, clustered FL generates multiple global models by clustering clients into several groups [13, 14, 15]. For example, an iterative clustering algorithm for FL named IFCA was proposed in [13], where client group identity estimation and federated model training are carried out alternatively. To estimate the group identity, clients are required to evaluate each global model on their local data to find the one achieving the minimal training loss. However, this process incurs additional training cost that increases linearly with the number of groups. More importantly, the clustering criterion in these works focuses on minimizing the training loss [13, 14, 27] but fails to reflect the generalization performance on the unseen data. Furthermore, selecting the number of groups is known to be challenging in these algorithms as the server does not know local data distribution.
@@ 3.1 Cross-Silo FL
Here $l(h(\mathbf{x};\mathbf{w}),y)$ is the loss function (e.g., the categorical cross-entropy) computed between the predicted label $h(\mathbf{x};\mathbf{w})$ and the ground-truth label $y$, and $\mathcal{P}_{i}^{te}$ denotes the distribution of test data. Although the test data are typically unknown during training, it is commonly assumed that the training data and test data are generated from the same underlying distribution [31], i.e.,
@@
In real-world scenarios, clients have various data distributions and different training dataset sizes [11, 32]. To be specific, some clients may own a limited number of data samples and thus seek assistance from others for collaboration to improve the model performance. To describe this scenario, without loss of generality, we assume that clients are partitioned into $K$ groups with indices $k\in[K]$. Note that the number of groups $K$ is a priori unknown. Each group consists of a set of clients $\mathcal{C}_{k}\subset\mathcal{N}$, and a client belongs to only one group. With the available training data $\hat{\mathcal{D}}_{k}^{tr}\triangleq\cup_{i\in\mathcal{C}_{k}}\mathcal{D}_{i}^{tr}$, clients in group $k$ train a model by minimizing the following objective:
@@
In training epoch $t$, the training process of each group is detailed as follows. Each client $i$ receives the global model $\hat{\mathbf{w}}_{G_{i}^{t}}^{t}$ from its corresponding group $G_{i}^{t}\in[K]$ and optimizes this model by using mini-batch stochastic gradient descent (SGD) for $Q$ steps. Specifically, at step $q\in[Q]$ of epoch $t$, client $i$ updates the local model according to
@@
$\mathbf{g}_{i}^{t}\triangleq\sum_{q=1}^{Q}\frac{1}{|\mathcal{B}_{i}^{t,q}|}\sum_{(\mathbf{x},y)\in\mathcal{B}_{i}^{t,q}}\nabla_{\mathbf{w}}l(h((\mathbf{x};\mathbf{w}_{i}^{t,q}),y),$ (5)
@@
It is worth noting that clients have various data distributions [1, 4], and thus some clients’ data may differ significantly from others. However, under strict privacy protection, the local data distribution is only visible to the client itself [33]. If the process of joining a group is performed arbitrarily, clients may suffer from performance degradation due to the data heterogeneity issue [11]. To clearly demonstrate this phenomenon, we present an example in the next subsection.
@@ 3.2 Motivating Example
We consider a cross-silo FL system with $N=3$ clients. They collaborate to classify data samples extracted from the CIFAR-10 [34] dataset by training a convolutional neural network (CNN) model with four convolutional layers. We assume that clients $1$ and $2$ have the same data distributions and client $3$ has different data distribution deviating from the two others. To be concrete, clients $1$ and $2$ have respectively $20\%$ and $80\%$ of five classes of training data, while client $3$ has all the training data samples of the other five classes. We simulate different collaboration patterns among three clients and evaluate their test errors, as shown in Table II, where the best generalization performances are highlighted in bold while the worst ones are underlined.
@@
Partial $(1,2)$ $\mathbf{23.54\pm 1.40}$ $\mathbf{25.28\pm 0.64}$ $\mathbf{15.90\pm 1.19}$ $\mathbf{21.58}$ $\mathbf{4.10}$
@@
Two clients collaborate on the global model, while the other client trains independently.
@@ 4.1 Theoretical Analysis of Generalization Performance
Consider the loss function with $\mu$-strong convexity and $L$-Lipschitz continuity. For a prediction function $h$ and any $10<\delta<$ 1, the following holds:
@@
with probability $1-\delta$, where $d_{1}\left(\hat{\mathcal{P}}_{G_{i}},\mathcal{P}_{i}\right)$ is the distribution divergence (e.g., $\mathcal{H}$-divergence) between $\hat{\mathcal{P}}_{G_{i}}$ and $\mathcal{P}_{i}$, and $\lambda=\min\left\{\mathbb{E}_{\mathcal{P}_{i}}\left[f_{G_{i}}(\mathbf{x})-f_{i}(\mathbf{x})\right],\mathbb{E}_{\hat{\mathcal{P}}_{G_{i}}}\left[f_{G_{i}}(\mathbf{x})-f_{i}(\mathbf{x})\right]\right\}$ with true labeling functions $f_{i}(\cdot)$ and $f_{G_{i}}(\cdot)$.
@@
The result is proved following [35, 36]. Please refer to Appendix A.
@@
The above results show that any client should seek collaborators with similar training data distributions and more data samples. However, the data distributions of clients cannot be disclosed to others due to privacy concerns. To solve this problem, in the following subsection, we propose a method to estimate the data divergence and formulate the client utility maximization problem.
4.2 Client Utility Maximization
@@
$s\left(i,G_{i}^{t}\right)=\cos\langle\mathbf{g}_{i}^{t},\hat{\mathbf{g}}_{{G_{i}^{t}}}^{t}\rangle=\frac{\left\langle\mathbf{g}_{i}^{t},\hat{\mathbf{g}}_{{G_{i}^{t}}}^{t}\right\rangle}{\left\|\mathbf{g}_{i}^{t}\right\|_{2}\left\|\hat{\mathbf{g}}_{{G_{i}^{t}}}^{t}\right\|_{2}}.$ (8)
@@
$U_{i}(G_{i}^{t})=U_{i}(\hat{D}_{G_{i}^{t}},\hat{\mathbf{g}}_{{G_{i}^{t}}}^{t})\triangleq-\frac{\alpha}{\hat{D}_{G_{i}^{t}}}+s\left(i,G_{i}^{t}\right)+\beta,$ (9)
where $\alpha>0$ is a trade-off constant, and $\beta>0$ is a large constant ensuring non-negative utility. Such non-negativity guarantees its physical meaning of utility without impacting system design. We note that $\alpha$ balances between training data volume and gradient similarity. If clients have sufficient data, they prefer collaborating with similar clients by adopting a small $\alpha$; otherwise, they are hungry for more data and choose a large $\alpha$. The impact of $\alpha$ is discussed further in Section 7.4 via simulations. According to the definition in (9), a client can improve its utility by involving more training data and increasing the similarity between gradients.
@@
Finding an optimal solution for Problem (P1) is $\mathcal{NP}$-hard in general [20]. One straightforward approach is to search over all $B_{N}$ possible partitions, where $B_{N}$ denotes the Bell number of a set with $N$ elements. This approach, however, introduces extremely high computational complexity. We notice that if $\alpha=0$, Problem (P1) reduces to a $k$-means problem which partitions $N$ clients into $K$ groups to minimize the within-cluster variance, i.e., the distance between each point $\mathbf{g}_{i}^{t}$ and its corresponding center $\mathbf{g}_{G_{i}}^{t}$. There are many classic methods for solving this problem, e.g., $k$-means clustering [39], $k$-means++ [40] and hierarchical clustering [41, 42] algorithms. However, the presence of the constraint $\alpha>0$ makes it unfeasible to define a suitable distance metric to solve Problem (P1). Consequently, these methods cannot be directly employed. Drawing inspiration from the aforementioned clustering algorithms, we introduce an algorithm to efficiently address the client partition in Problem (P1) in the following section.
@@ 5 Hierarchical Clustering-based Collaborative Training Scheme
6:        Partition clients according to $\{\mathcal{C}_{k}\}_{k=1}^{K}\leftarrow\textit{ClientPartition}(\{\mathbf{g}_{i}^{t}\}_{i\in\mathcal{N}})$;
@@
where $\hat{D}_{k^{\prime}}=\sum_{j\in\mathcal{C}_{k^{\prime}}}D_{j}$ is the number of training data samples of clients in $\mathcal{C}_{k^{\prime}}$. For these clients in group $k^{\prime}$, we recompute their utilities and sum them up as the utility of this group, i.e.,
@@
It is worth noting that there is no need to select beforehand the number of groups $K$. In every partitioning step, the value of $K$ will decrease by one automatically. Thus, the number of groups is implicitly determined by the stopping criterion. Specifically, the process of client partition is terminated if there is only one group left, i.e., $K=1$, or there is no benefit in forming collaborations between any two groups, i.e.,
@@
Through multiple steps, HCCT builds a hierarchical representation of clients and finalizes the grouping among them. The process of client partition is summarized in Algorithm 2, and we show an example with six clients in Fig. 2.